# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

### v1

@@ 4.4 Sonnet with different attention variants
To understand how different attention mechanisms affect the forecasting accuracy of Sonnet, we replace the MVCA module with the aforementioned modified attention mechanisms (see section 4.3). Experiments are conducted on both the ILI and WEA forecasting tasks across all locations, forecasting horizons, and test periods. Hyperparameters are re-tuned for each attention variant using the same validation procedure. Further experimental details are provided in Appendix C.2.
Table 3: Performance of Sonnet across the ILI and WEA tasks with different attention modules. Best results are bolded and second best are underlined.
Sonnet VDAB [43] FED [62] FNet [21] Naïve
ILI-ENG 7 1.4791 21.84 1.4327 20.33 5.9913 77.67 1.5908 27.67 1.7170 25.34
14 1.9225 25.77 1.9729 29.39 6.0309 78.05 2.0841 32.81 2.0677 31.42
21 2.5101 36.53 2.7744 35.16 6.0447 78.24 2.7652 39.31 2.7812 37.17
28 2.7481 36.95 3.0602 44.31 6.0508 78.38 3.2403 44.51 3.0785 50.03
ILI-US2 7 0.3806 14.89 0.4294 16.57 1.1675 45.00 0.4146 18.50 0.4356 17.55
14 0.4491 18.38 0.4997 19.10 1.1792 45.53 0.5585 23.15 0.5316 20.43
21 0.5326 20.64 0.5748 20.49 1.1954 46.19 0.6593 25.31 0.6130 23.46
28 0.5788 21.15 0.6426 24.26 1.1934 46.13 0.7053 29.09 0.6514 26.12
ILI-US9 7 0.2668 12.98 0.2766 13.37 0.8558 40.97 0.2857 14.44 0.3006 13.85
14 0.2806 13.10 0.3239 15.81 0.8560 40.98 0.3510 16.13 0.3398 14.89
21 0.3179 14.11 0.3188 14.50 0.8567 41.01 0.3788 17.79 0.4038 19.23
28 0.3675 16.53 0.3748 16.81 0.8570 41.02 0.3791 18.20 0.4111 19.07
WEA-CT 4 1.6240 9.63 1.7135 10.11 4.3594 24.65 1.6287 9.63 1.7260 10.21
12 3.5432 20.38 3.5631 20.48 4.3595 24.65 3.5880 20.56 3.5581 20.47
28 3.7277 21.32 3.7602 21.50 4.3585 24.65 3.9155 22.32 3.7584 21.48
120 3.7373 21.35 3.7339 21.33 4.3564 24.64 3.8150 21.73 3.7632 21.50
Sonnet VDAB [43] FED [62] FNet [21] Naïve
WEA-HK 4 0.6389 4.05 0.6451 4.11 2.8011 17.74 0.6389 4.08 0.6760 4.30
12 1.2355 7.74 1.2786 8.01 2.8008 17.74 1.3254 8.33 1.2810 8.06
28 1.4135 8.83 1.4633 9.06 2.8078 17.78 1.4941 9.30 1.4745 9.17
120 1.5469 9.58 1.5469 9.60 2.8100 17.80 1.6389 10.15 1.6211 10.04
WEA-LD 4 1.7231 15.16 1.8135 15.79 4.6682 34.93 1.7402 15.30 1.8107 15.99
12 2.9589 23.88 2.9863 24.06 4.6614 34.90 3.0078 24.11 2.9733 23.95
28 3.2161 25.49 3.2712 25.85 4.6593 34.87 3.3784 26.45 3.2682 25.84
120 3.2464 25.82 3.3193 26.24 4.6676 34.91 3.6666 28.48 3.4072 26.73
WEA-NY 4 1.2716 11.94 1.3307 12.36 5.3865 40.92 1.2684 11.77 1.3661 12.92
12 2.4476 21.23 2.4742 21.56 5.3836 40.91 2.5101 21.80 2.4380 21.45
28 2.6744 23.10 2.7081 23.43 5.3871 40.93 2.7972 23.81 2.7238 23.11
120 2.7135 23.15 2.7876 23.55 5.3892 40.93 3.0629 25.58 2.9106 24.55
WEA-SG 4 0.3444 1.25 0.3461 1.26 0.6999 2.54 0.3409 1.24 0.3455 1.26
12 0.4160 1.51 0.4207 1.53 0.6976 2.53 0.4238 1.54 0.4246 1.54
28 0.4653 1.69 0.4729 1.72 0.6979 2.53 0.4947 1.80 0.4712 1.71
120 0.4830 1.76 0.4957 1.80 0.6974 2.53 0.5119 1.86 0.4850 1.76
@@
Results are enumerated in Table 3. Overall, using MVCA yields the best accuracy compared to other attention modules, outperforming them in $26$ out of the $32$ tasks (MAE). On average, it reduces the MAE and sMAPE scores by $2.9\%$ and $2\%$, respectively. The most competitive baseline module is VDAB (DeformTime). MVCA outperforms it by $3.6\%$ and $3\%$ on average w.r.t. MAE and sMAPE, respectively. The worst-performing attention module is FED (FEDformer). MVCA reduces its MAE by $51.5\%$. FED focuses on capturing seasonality information using frequency transformation without considering the dependencies between variables. This shows that modelling seasonality alone is insufficient for MTS forecasting, particularly when key predictive information comes from exogenous variables. The performance obtained by using either naïve attention or FNet is similar. MVCA reduces their MAE by $7\%$ and $6.6\%$ on average, respectively. This is an expected outcome as FNet does not introduce learnable parameters to the naïve attention.
Table 4: Ablation study results across the ILI (4 test seasons) and WEA tasks (3 test years). MAE is shown for each ablated variation ($\neg$ denotes removal of a component). $\Delta\%$ is the MAE percentage of difference compared to Sonnet’s MAE (which can be seen in Table 1).
$\neg$ Coherence $\neg$ MLP $\neg$ MVCA $\neg$ Embed $\neg$ Koopman
$H$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$
ILI-ENG 7 1.4224 –3.83 1.5312 +3.52 1.5906 +7.54 1.6333 +10.42 1.4855 +0.43
14 2.1099 +9.75 2.1710 +12.92 2.2346 +16.23 2.0658 +7.45 2.0793 +8.15
21 2.8797 +14.72 2.6141 +4.15 2.7939 +11.31 2.6467 +5.44 2.6299 +4.77
28 3.4200 +24.45 3.1145 +13.33 3.2745 +19.15 3.0910 +12.47 3.0622 +11.43
ILI-US2 7 0.4142 +8.82 0.4109 +7.96 0.4317 +13.43 0.4162 +9.37 0.4205 +10.48
14 0.5120 +14.00 0.5074 +12.96 0.5277 +17.50 0.5295 +17.90 0.4963 +10.51
21 0.6087 +14.29 0.5963 +11.96 0.6711 +26.02 0.5729 +7.57 0.5901 +10.81
28 0.6233 +7.67 0.6622 +14.39 0.6659 +15.05 0.6065 +4.78 0.6338 +9.50
ILI-US9 7 0.2720 +1.93 0.3017 +13.10 0.2775 +4.01 0.2784 +4.36 0.2796 +4.80
14 0.3183 +13.41 0.3341 +19.05 0.3307 +17.85 0.3146 +12.12 0.3151 +12.30
21 0.3575 +12.47 0.3264 +2.68 0.3929 +23.60 0.3113 –2.08 0.3192 +0.41
28 0.4277 +16.40 0.3446 –6.23 0.4013 +9.20 0.3587 –2.39 0.3476 –5.40
WEA-CT 4 1.6633 +2.42 1.6434 +1.20 1.6355 +0.71 1.6779 +3.32 1.6495 +1.57
12 3.5909 +1.35 3.5743 +0.88 3.5886 +1.28 3.5722 +0.82 3.5677 +0.69
28 3.7832 +1.49 3.7605 +0.88 3.8620 +3.60 3.7654 +1.01 3.7543 +0.72
120 3.7695 +0.86 3.7639 +0.71 3.7685 +0.83 3.7662 +0.77 3.7667 +0.79
$\neg$ Coherence $\neg$ MLP $\neg$ MVCA $\neg$ Embed $\neg$ Koopman
$H$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$
WEA-HK 4 0.6390 +0.00 0.6395 +0.08 0.6310 –1.24 0.6553 +2.55 0.6394 +0.07
12 1.2904 +4.45 1.2607 +2.04 1.3023 +5.41 1.2779 +3.43 1.2733 +3.07
28 1.4747 +4.33 1.4671 +3.80 1.4770 +4.50 1.4667 +3.76 1.4626 +3.48
120 1.6347 +5.68 1.6267 +5.16 1.6199 +4.72 1.5991 +3.37 1.6039 +3.68
WEA-LD 4 1.7538 +1.78 1.7675 +2.58 1.7420 +1.10 1.7868 +3.70 1.7555 +1.88
12 2.9906 +1.07 3.0070 +1.63 3.0009 +1.42 3.0187 +2.02 2.9926 +1.14
28 3.3126 +3.00 3.2812 +2.02 3.3278 +3.47 3.3261 +3.42 3.2681 +1.62
120 3.3854 +4.28 3.3468 +3.09 3.3749 +3.96 3.3591 +3.47 3.3454 +3.05
WEA-NY 4 1.2962 +1.94 1.2941 +1.77 1.2806 +0.71 1.3158 +3.48 1.2841 +0.99
12 2.4466 –0.04 2.4643 +0.68 2.4685 +0.85 2.4666 +0.78 2.4595 +0.49
28 2.7515 +2.88 2.7084 +1.27 2.7644 +3.36 2.7392 +2.42 2.7230 +1.82
120 2.8564 +5.27 2.7774 +2.36 2.9380 +8.27 2.8401 +4.67 2.8013 +3.24
WEA-SG 4 0.3417 –0.79 0.3446 +0.05 0.3426 –0.52 0.3481 +1.09 0.3431 –0.36
12 0.4226 +1.60 0.4194 +0.81 0.4260 +2.41 0.4228 +1.62 0.4220 +1.43
28 0.4894 +5.18 0.4755 +2.20 0.4742 +1.91 0.4754 +2.17 0.4759 +2.28
120 0.4999 +3.49 0.4942 +2.31 0.5028 +4.10 0.4965 +2.80 0.4950 +2.49
4.5 Ablation study on Sonnet
An ablation study is performed to better understand the contribution of Sonnet’s components to the overall forecasting accuracy. We ablate over the following modules: the coherence mechanism (within MVCA), the MLP head for inter-variable dependency modelling (within MVCA), the complete MVCA module (section 3.3), the joint embedding module (section 3.1), and the Koopman evolvement module (section 3.4). A more detailed description of each ablation variant is provided in Appendix D.9. We re-tune Sonnet’s hyperparameters for each ablation variant across all locations, test seasons, and forecasting horizons (for the ILI and WEA tasks).
Results are enumerated in Table 4. We report the average MAE across all test seasons and its percentage of difference compared to the Sonnet model (enumerated in Table 1). Among all the ablated components, the MVCA module has the greatest impact. Removing MVCA results in an increase of MAE by $6.3\%$ on average across the ILI and WEA tasks. Breaking down MVCA further, we observe that removing its two internal components—coherency modelling and the MLP residual connection—leads to notable drops in performance ($5.2\%$ and $4.1\%$ increase in MAE, respectively), but less pronounced than removing the entire MVCA block. This indicates that while capturing coherence over time and enhancing variable interactions with residual connected MLP are both important modules, they also work in tandem within the full MVCA design.
@@ Appendix A Data sets for multivariable time series forecasting
In this section, we provide a detailed description of all the data sets used in our MTS forecasting experiments. The weather data set that is formed by us is available for download at github.com/ClaudiaShu/Sonnet.
@@ A.1 Hourly electricity transformer temperature prediction
The Hourly Electricity transformer temperature (ETTh1 and ETTh2) data sets are two widely used benchmarks in time series forecasting, obtained from $2$ different counties in China [61]. [^2] ETTh1 and ETTh2, github.com/zhouhaoyi/ETDataset The target variable of both data sets is the oil temperature. They contain $6$ exogenous variables capturing power load attributes. Both data sets cover the period from July 1, 2016, to June 26, 2018. We use only a subset of $14{,}400$ time steps starting from July 1, 2016, to match prior evaluation setups of the models that we are comparing to [34, 25, 43]. Within the subset, the first $8{,}640$ time steps are used for training, the next $2{,}880$ for validation, and the last $2{,}880$ for testing.
@@ A.2 Electricity consumption prediction
The Electricity Consumption (ELEC) data set provides the electricity consumption records for Zurich (Switzerland), collected every $15$ minutes. [^3] Electricity consumption data set, github.com/unit8co/darts ELEC contains measurements for both residential households (recorded at low voltage) and commercial businesses (recorded at medium voltage). We choose the electricity consumption of households as our target variable for prediction. External predictors of the electricity consumption include $8$ weather-related covariates recorded hourly by a weather station in the city of Zurich.
Note that for some time intervals, certain predictor values may be missing. These are imputed using linear interpolation. We then resample rates to an hourly frequency following prior work in electricity consumption prediction [1, 57].
For the electricity consumption task, we use two test periods, years 2020 ($8{,}784$ time steps, including the leap day) and 2021 ($8{,}760$ time steps). The year preceding the test year is used as the validation ($8{,}760$ and $8{,}784$ time steps, respectively), while the $4$ years before validation form the training set ($35{,}064$ in both evaluations).
@@ A.3 Energy price prediction
The energy price (ENER) data set contains $4$ years of electrical consumption, generation, and pricing data, with the main target of predicting the actual energy price in Spain. [^4] Energy price data set, github.com/unit8co/darts Note that the original data sets have $27$ exogenous variables in total. However, $8$ variables consist entirely of zeros or missing values, which do not provide useful information for forecasting purposes. We therefore exclude these variables from the analysis and use the remaining $19$. Data is sampled on an hourly basis. Data from December 31, 2014 to December 31, 2016 are used for training ($17{,}545$ samples). The subsequent year (January 1 to December 31, 2017, $8{,}760$ samples) is used for validation. For testing, we use data from January 1, 2018, to 22:00, December 31, 2018 ($8{,}759$ samples); the price for the last hour is not available.
A.4 Weather forecasting on selected locations
We form the weather forecasting (WEA) data set using climate information extracted from the WeatherBench repository [38], covering the period from 1979 to 2018.555WeatherBench repository, dataserv.ub.tum.de/index.php/s/m1524895 Data is sampled from a uniform spatial grid over the globe with a resolution of $5.625^{\circ}$ in both latitude and longitude. Each grid point represents a distinct geographical location. To capture different climatic conditions, we select five spatially diverse cities from the grid: London (UK), New York (US), Hong Kong (China), Cape Town (South Africa), and Singapore. Among these, London, New York, and Hong Kong are located in the northern hemisphere, with Hong Kong being in the subtropical zone; Singapore lies near the equator, and Cape Town is in the southern hemisphere. The two cities that are closer to the equator (Hong Kong and Singapore) do not have strong seasonal temperature fluctuations over the year. Consequently, we expect models to yield lower prediction errors at these two locations. For each city, climate data is sampled from the nearest grid point to its coordinates. [^6] We obtain the longitude and latitude of each city using OpenStreetMap (wiki.openstreetmap.org). Specifically, we obtain the absolute distance in both latitude and longitude between each city and all grid points, and select the grid point with the smallest distance along both axes. The forecasting target is the $850$ hPa temperature in Kelvin (T850) at the central grid point, [^7] Further information about T850 is provided at charts.ecmwf.int/products/medium-z500-t850. which is an important climate indicator that has been widely used in climate research [41, 14]. To form exogenous inputs, we include four other climate variables at each selected grid point following prior work [33, 49]. These are the $500$ hPa geopotential (Z500), $2$-meter air temperature (t2m), $10$-meter zonal wind (u10), and $10$-meter meridional wind (v10). We also include these $4$ variables and T850 from the surrounding eight grid points, forming a $3\times 3$ spatial window centred on each selected grid, resulting in $44$ exogenous variables in total. The data is resampled to maintain a temporal resolution of $6$ hours similarly to Verma et al. [49].
We consider $3$ test seasons: years 2016, 2017, and 2018 (Jan. 1 to Dec. 31). Each test year contains $1{,}464$, $1{,}460$, and $1{,}460$ time steps, respectively. The year immediately preceding a test year is used for validation (2015, 2016, and 2017). Validation year 2015 contains $1{,}460$ time steps. The training sets include data from January 1, 1980, up to the date preceding the corresponding validation year, resulting in $51{,}136$, $52{,}596$, and $54{,}060$ time steps for the three test seasons.
A.5 Influenza-like-illness rate prediction
We construct an ILI forecasting data set for three regions: England (UK), and US HHS Regions 2 and 9, following the setup in prior work [43]. Ground truth ILI rates are obtained from the Royal College of General Practitioners (RCGP) in England and the Centers for Disease Control and Prevention (CDC) for the US regions. Weekly ILI rates are linearly interpolated to a daily resolution. Specifically, we assume the ILI rates are reported corresponding to the middle day of each week, which is Wednesday for the US regions and Thursday for England [43]. We also incorporate the ILI rate reporting delay of syndromic surveillance systems; in England there is a lag of $\delta=7$ days and $\delta=14$ days for the US regions [50, 39].
We use web search activity data (daily frequency time series of different search terms), obtained from the Google Health Trends API, as the exogenous indicators. We follow the setup of DeformTime [43] to form the set of search queries for our experiments. [^8] Search queries can be obtained from github.com/ClaudiaShu/DeformTime. Note that for the US regions, search frequencies are collected at the state level, and the query data for each region is obtained using weighted averages based on the population of the states inside each region (New Jersey/New York is $0.32/0.68$ for Region 2 and Arizona/California is $0.16/0.84$ for Region 9).
We evaluate forecasting accuracy using $4$ test seasons (2015/16, 2016/17, 2017/18, and 2018/19) in England and the US regions, each corresponding to a 12-month period that encompasses at least one seasonal flu outbreak. For England, each season begins on September $1$ and ends on August $31$ of the next year, while for the US regions, it is from August $1$ to July $31$ of the following year. For each test season, models are trained on the preceding $9$ seasons. Validation sets are extracted from the last three seasons in the training data. Each validation set covers $180$ days, containing three $60$-day segments that capture the onset (from the last season), peak (from the penultimate season), and outset (the $3$rd to last season) phases of an influenza season. The onset is defined as the first point followed by two weeks above a predefined ILI threshold; the peak is the point of the highest ILI rate; and the outset is the last point followed by two weeks above the threshold. These selected points are then considered as the $30$th day in a $60$-day validation period.
Appendix B Further information about MVCA
In Equation 2, the cross-spectral and power-spectral density matrices are obtained using the frequency domain embeddings $\mathbf{Q}_{f}$, $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$. The obtained matrices $\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}$ and $\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}$ are both containing real numbers because multiplying a complex number (e.g. $z=a+ib$) by its conjugate ($a-ib$) results in its squared magnitude ($a^{2}+b^{2}$), which is a real number.
Appendix C Baseline forecasting models and attention mechanisms
C.1 Baseline MTS forecasting models
DeformTime [43] is specifically designed for multivariable forecasting. It uses deformable attention to capture the dependencies across time and covariates, reaching competitive performance on MTS forecasting tasks. Based on our evaluation, DeformTime is the best-performing baseline.
ModernTCN [27] applies convolution along the temporal and variable axes. By leveraging large receptive fields, it effectively captures both intra-variable and inter-variable dependencies.
TimeXer [52] is a transformer-based model that focuses on multivariable forecasting. It first embeds the exogenous variables along the sequence dimension, and then captures the exogenous information through a cross-attention mechanism.
Samformer [18] first embeds the input along the sequence dimension and then applies naïve attention along the feature dimension. It conducts forecasts using a linear layer. We also include Samformer as one of the base models that deploy naïve attention.
iTransformer [25] embeds the input along the sequence dimension and deploys multiple attention layers along the variable dimension. We also include iTransformer as one of the base models.
PatchTST [34] segments time series data into patches of fixed size, treating them as input tokens to a transformer. It processes each variable independently, enabling efficient modelling of long-term dependencies. PatchTST is also used as one of the base models.
Crossformer [60] captures both inter-variable and intra-variable dependencies using a two-stage attention mechanism, with each stage dedicated to one type of dependency. The information is captured using time series patches.
DLinear [58] uses a simpler neural network architecture (feedforward layers). It decomposes each input series into trend and seasonal components, which are then forecasted independently using linear layers. It models each variable separately, using its own historical values for prediction. Based on our evaluation, DLinear is the worst-performing model in MTS forecasting. It does not outperform persistence in the majority of our experiments (see Table 1). This corroborates findings in related work [43].
Naïve forecasting models. We include simple baseline models as reference points for performance comparison. We use either seasonal persistence or standard persistence, depending on whether a target variable (and task) is characterised by strong seasonality or not. A standard persistence model is a naïve baseline that uses the last observed value of the target variable as the prediction (forecasting outcome). The seasonal persistence model uses the value from the same point in the previous seasonal cycle. In our case, the electricity consumption data (ELEC) has a strong seasonal trend, with seasonal persistence outperforming the standard one by a large margin (results not shown). Therefore, we use the seasonal persistence model for the ELEC task, while for the other tasks, we retain standard persistence.
Table S1: ILI rate forecasting results in England (ILI-ENG) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.

### v2

@@ 4.4 Sonnet with Different Attention Variants
To understand how different attention mechanisms affect the forecasting accuracy of Sonnet, we replace the MVCA module with the aforementioned modified attention mechanisms (see section 4.3). Experiments are conducted on both the ILI and WEA forecasting tasks across all locations, forecasting horizons, and test periods. Hyperparameters are re-tuned for each attention variant using the same validation procedure. Appendix C.3 provides further details.
@@
Results are enumerated in Table 3. Overall, using MVCA yields the best accuracy compared to other attention modules, outperforming them in $26$ out of the $32$ tasks (MAE). On average, it reduces the MAE and sMAPE scores by $2.9\%$ and $2\%$, respectively. The most competitive baseline module is VDAB (DeformTime). MVCA outperforms it by $3.6\%$ and $3\%$ on average w.r.t. MAE and sMAPE, respectively. The worst-performing attention module is FED (FEDformer). MVCA reduces its MAE by $51.5\%$. FED focuses on capturing seasonality information using frequency transformation without considering the dependencies between variables. This shows that modelling seasonality alone is insufficient for MTS forecasting, particularly when key predictive information comes from exogenous variables. The performance obtained by using either naïve attention or FNet is similar. MVCA reduces their MAE by $7\%$ and $6.6\%$ on average, respectively. This is expected as FNet does not introduce learnable parameters to the naïve attention.
@@ Appendix A Data sets for multivariable time series forecasting
In this section, we provide a detailed description of all the data sets used in our MTS forecasting experiments. The weather data set that is formed by us is available in our code and data repository.
@@ A.1 Hourly electricity transformer temperature prediction
The Hourly Electricity transformer temperature (ETTh1 and ETTh2) data sets are two widely used benchmarks in time series forecasting, obtained from $2$ different counties in China (zhou2021informer). [^3] ETTh1/h2, github.com/zhouhaoyi/ETDataset The target variable of both data sets is the oil temperature. They contain $6$ exogenous variables capturing power load attributes. Both data sets cover the period from July 1, 2016, to June 26, 2018. We use only a subset of 14,40014{,}400 time steps starting from July 1, 2016, to match prior evaluation setups of the models that we are comparing to (nie2023time; liu2024itransformer; shu2025deformtime). Within the subset, the first 8,6408{,}640 time steps are used for training, the next 2,8802{,}880 for validation, and the last 2,8802{,}880 for testing.
@@ A.2 Electricity consumption prediction
The Electricity Consumption (ELEC) data set provides the electricity consumption records for Zurich (Switzerland), collected every $15$ minutes. [^4] Electricity data set, github.com/unit8co/darts ELEC contains measurements for both residential households (recorded at low voltage) and commercial businesses (recorded at medium voltage). We choose the electricity consumption of households as our target variable for prediction. External predictors of the electricity consumption include $8$ weather-related covariates recorded hourly by a weather station in the city of Zurich.
Note that for some time intervals, certain predictor values may be missing. These are imputed using linear interpolation. We then resample rates to an hourly frequency following prior work in electricity consumption prediction (møllerandersen2017households; yukseltan2020hourly).
For the electricity consumption task, we use two test periods, years 2020 (8,7848{,}784 time steps, including the leap day) and 2021 (8,7608{,}760 time steps). The year preceding the test year is used as the validation (8,7608{,}760 and 8,7848{,}784 time steps, respectively), while the $4$ years before validation form the training set (35,06435{,}064 in both evaluations).
@@ A.3 Energy price prediction
The energy price (ENER) data set contains $4$ years of electrical consumption, generation, and pricing data, with the main target of predicting the actual energy price in Spain. [^5] Energy price data set, github.com/unit8co/darts Note that the original data sets have $27$ exogenous variables in total. However, $8$ variables consist entirely of zeros or missing values, which do not provide useful information for forecasting purposes. We therefore exclude these variables from the analysis and use the remaining $19$. Data is sampled on an hourly basis. Data from December 31, 2014 to December 31, 2016 are used for training (17,54517{,}545 samples). The subsequent year (January 1 to December 31, 2017; 8,7608{,}760 samples) is used for validation. For testing, we use data from January 1, 2018, to 22:00, December 31, 2018 (8,7598{,}759 samples); the price for the last hour is not available.